# Quantum correlations under decoherence

### v1

@@ III The model, the master equation and its solution
Our model system is a Heisenberg chain with only two localised qubits having the spin-spin interaction $J_{x}S_{x}^{(1)}S_{x}^{(2)}+J_{y}S_{y}^{(1)}S_{y}^{(2)}+J_{z}S_{z}^{(1)}S_{z}^{(2)}$ and in interaction with an applied magnetic field along $z$-direction. The Hamiltonian for the $XY(J_{z}=0)$ model then recasts
$H$ $=$ $J(S_{+}^{(1)}\otimes S_{-}^{(2)}+S_{-}^{(1)}\otimes S_{+}^{(2)})+\Delta(S_{+}^{(1)}\otimes S_{+}^{(2)}+S_{-}^{(1)}\otimes S_{-}^{(2)})$ (43)
@@
where $S_{\pm}^{(i)}=S_{x}^{(i)}\pm iS_{y}^{(i)}$ are the spin raising (lowering) operators for the $i^{th}$ qubit and $\omega$ is the intensity of the applied magnetic field. Coupling between qubits have been denoted as $J=(J_{x}+J_{y})/4$ and $\Delta=(J_{x}-J_{y})/4$ where the latter quantifies the anisotropicity of the interaction. It is noteworthy that systems with two coupled qubits have been utilized extensively in various contexts such as cavity QED GyMeAw-PRB-2006 and nuclear spin systems with the hope of possible applications in quantum information schemes and quantum computing NiCh-book-2000 ; MiLaSaKn-PRA-2002 .
At zero temperature, $T=0$, the state of the system evolves under the Lindbladian master equation Waetal-JPB-2006
$\dot{\rho}(t)$ $=$ $-i[H,\rho(t)]+\gamma~{}\mathcal{D}[S_{-}^{(1)}\otimes\mathds{1}^{(2)}]\rho(t)+\gamma~{}\mathcal{D}[\mathds{1}^{(1)}\otimes S_{-}^{(2)}]\rho(t),$ (44)
where $\mathcal{D}[A]B=ABA^{\dagger}-\{A^{\dagger}A,B\}/2$ is the Lindblad super-operator and $\{X,Y\}$ denotes the anti-commutator of operators $X$ and $Y$. In our open quantum system this Lindblad equation describes the population relaxation of the upper state of each qubit with the relaxation rate $\gamma$ which for simplicity it has been taken equal for both qubits. In other words, interaction of qubits with the environment is assumed to be the same. As discussed in Waetal-JPB-2006 , this assumption is reasonable when the rotating wave approximation remains valid. The Lindblad equation (44) maintains the shape X of the state i.e., if the initial state is a X-shaped i.e., of the form (6) then the evolved state $\rho(t)$ is too.
@@
where $|\Psi_{-}\rangle=\frac{1}{\sqrt{2}}(|01\rangle-|10\rangle)$ is one of the maximally entangled Bell states, and $p$ a real number with constraint $p\in[-\frac{1}{3},1]$ which comes from the non-negativity requirement of the density matrix. For $p\in[-\frac{1}{3},\frac{1}{3}]$ the Werner state is separable and for $p\in[-\frac{1}{3},\frac{1}{\sqrt{2}}]$ it is local Luc-thesis-2019 . From this one sees that the Werner state is entangled but local for $p\in(\frac{1}{3},\frac{1}{\sqrt{2}}]$ i.e., entanglement does not imply nonlocality. The purity of the Werner state (62) is $\text{tr}(\rho_{W}^{2})=(1+3p^{2})/4$ being one for $p=1$ and $1/4$ for $p=0$. The former implies the pure ensemble $|\Psi_{-}\rangle\langle\Psi_{-}|$ while the latter corresponds to the completely random ensemble. The solution of the Lindblad equation (44) this time is given by
@@
where $\Omega$ is given by (57). As one sees the state $\rho_{W}(t)$ is independent of the parameter $J$ appearing in the Hamiltonian being responsible for the isotropic part of the spin-spin interaction. An interesting point about the Werner state is that for $\gamma=0$, density operator is independent of time which is the manifestation of the fact that the Werner state (62) commutes with the Hamiltonian (43) of the system, $[H,\rho_{W}(t)]=0$.
@@
for the steady state to be entangled. As we saw, reduced density operators for a X-shaped state are both always diagonal with zero quantum coherence. However, Eq. (74) shows that the global steady state $\ell_{1}$-norm coherence is
@@ IV Temperature dependence of the steady state solution
$\dot{\rho}(t)$ $=$ $-i[H,\rho(t)]+\gamma(\bar{n}+1)~{}\mathcal{D}[S_{-}^{(1)}\otimes\mathds{1}^{(2)}]\rho(t)+\gamma(\bar{n}+1)~{}\mathcal{D}[\mathds{1}^{(1)}\otimes S_{-}^{(2)}]\rho(t)$ (77)
$+\gamma\bar{n}~{}\mathcal{D}[S_{+}^{(1)}\otimes\mathds{1}^{(2)}]\rho(t)+\gamma\bar{n}~{}\mathcal{D}[\mathds{1}^{(1)}\otimes S_{+}^{(2)}]\rho(t),$
$\bar{n}$ being the average excitation of the thermal bath parametrizing the temperature: being zero for $T=0$ Waetal-JPB-2006 . This equation is known as quantum optical master equation where $\gamma$ is the rate for spontaneous emission and $\gamma\bar{n}$ describes the rate of thermally induced emission and absorption processes BrPe-book-2002 222page: 148, equation (3.219). Eq. (77) can be analytically solved for the X-state (6). Again, one sees that this master equation preserves the shape of the state. Here, we are only interested in the steady state, $t\to\infty$, solution which is given by
@@
respectively for the steady state concurrence, CC and MIN. In this case the $W$ matrix (18) is diagonal with $W_{11}(\infty)=W_{22}(\infty)$. Thus, for the steady state LQU one has that
@@ V Results and discussion
In this section, we aim to consider the impact of the relaxation rate $\gamma$, responsible for the interaction of our qubit-qubit Heisenberg XY system with the environment, on the dynamics of quantum correlations by fixing other parameters of the problem including coupling strengths $J$ and $\Delta$; and magnetic field intensity $\omega$. However, as wee see in the previous section, Werner states and thus corresponding quantum correlations are independent of $J$. The same is valid for the general X-shaped state in the steady situation. Furthermore, in some calculations we fix all parameters other than $\Delta$ to study anisotropicity dependence of quantum correlations. Our investigation will primarily focus on examining how the zero-temperature steady state quantum correlations appear for different initial X states, including a mixture and the Werner state. After, the interesting phenomenon of entanglement sudden death, for independent qubits, will be considered for two initially entangled and afterwards not interacting qubits both for zero and finite temperatures. The time of ESD will be obtained analytically (numerically) for zero (finite) temperatures. Environment-induced quantum correlations is another phenomenon which will be considered. At the end finite temperature quantum correlations will be examined in the steady state case by studying their behaviour with temperature.
@@
Figure 1 depicts evolution of different quantum correlations for $\omega=1$, $J=0.1$ and $\Delta=0.5$ for three different values of the relaxation rate $\gamma=0.1$, $\gamma=0.15$ and $\gamma=0.2$. Here, system is initially described by the entangled mixture (49). Initial values of the quantum correlations, being independent of $\gamma$, are $\mathcal{C}(0)=0.5$, $\mathcal{L_{N}}(0)=0.2716$, $\mathcal{U}(0)=0.5$ and $C_{cc}(0)=0.5$. As this figure reveals, both entanglement quantifiers, concurrence and log negativity, exhibit several dark and revival periods FiTa-PRA-2006 before reaching the steady state. Because of the spontaneous emission, the initial entanglement diminishes with time, disappears in a given finite time which itself depends of the value of $\gamma$ and after a dark while it revives. The number and length of these dark intervals both depend on the value of $\gamma$. The first dark period is the longest one which itself decreases with $\gamma$. The difference between entanglement quantifiers, $\mathcal{C}$ and $\mathcal{L_{N}}$, and quantum correlations beyond entanglement, $\mathcal{U}$ and $C_{cc}$, is that the former ones revive after zero entanglement, while the others can revive immediately several times without completely vanishing for a period of time; there are no dark periods for quantum correlations beyond QE. Moreover, LQU captures less quantum correlations than the other quantifiers, with the former converging to a stationary value of $0.1597$ for $\gamma=0.1$. As one expects, the transition to the steady state occurs sooner for higher values of $\gamma$ and as one can check analytically from Eq. (74) the steady state value of quantum correlation although is independent of the initial state but depends on the value of $\gamma$.
@@
Figure 1: Evolution of the quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel), for $\gamma=0.1$ (red curves), $\gamma=0.15$ (green curves), $\gamma=0.2$ (blue curves) for $\omega=1$, $J=0.1$ and $\Delta=0.5$. The quantum system is initially described by the mixture (49).
@@
Then one has that $1-\text{tr}(\rho_{\text{mix}}^{2})=2w(1-w)$ for the purity of the state. Time dependent density matrix obtained from the solution of the Lindblad equation (44) with $J=\Delta=0$ and the initial condition (91) is then,
@@
From the definition (8) one obtains
@@
for the concurrence of this mixture. This equation shows that this state is always separable for $w=1$ as it is obvious from (91) for the initial state. For $10\leq w<$ 1 which the initial state is entangled, Eq. (99) shows that there is a finite time where the entanglement disappears, the signature of ESD. This time is obtained as
@@
which is infinite for $w=0$ meaning that the entanglement does not die suddenly for the maximally entangled state $\frac{1}{2}(|00\rangle+|11\rangle)(\langle 00|+\langle 11|)$. The age of entanglement is shorter for higher values of both $\gamma$ and $w$. See figure 2.
Figure 2: Left panel: time of entanglement death $\gamma\tau$ given by Eq. (100) in terms of the weight $w$ appearing in the mixture (91). Right panel: evolution of concurrence (99) for $w=0.1$ (black curve), $w=0.3$ (red curve), $w=0.5$ (green curve) and $w=0.8$ (blue curve). Parameters have been chosen $\omega=1$ and $\Delta=J=0$ i.e, qubits solely interact with the common environment not with each other.
@@
for CC or equivalently for MIN with respect to (30). This equation shows that $C_{cc}(0)=1-w$, being nonzero for $w\neq 1$, and CC decreases exponentially; the usual behaviour of decaying systems. Thus, there is no phenomenon of CC sudden death i.e., this quantity does not disappear in a finite time. For LQU which is computed by (23) analytical calculations show that the square root expression appearing in this equation is zero. Thus, LQU for $J=\Delta=0$ is obtained from
@@
where $W_{ij}$ is given by (18). The initial value of LQU is $\mathcal{U}(0)=\max\{1-w,1-\sqrt{w(1-w)}\}$ being zero for $w=1$. In the left panel of Fig. 3 we have plotted initial LQU versus the probability weight. As the numerical computations in the right panel of Fig. 3 show, LQU decreases first reaching a local minimum then increases to a given value and breaks downwards and continuously, not abruptly, vanishes. Ref. GuPeZeWa-QIP-2020 has already reported this sudden change in LQU when one of the qubits is subjected to local decoherence noise. In any way, there is no LQU sudden death. Note that breaking points in the curves are due to the “max” operation in the definition (102).
@@
Figure 3: Initial value of LQU in terms of the probability weight $w$ (left panel) and evolution of LQU (right panel) for $w=0.4$ (cyan), $w=0.5$ (red), $w=0.6$ (green) and $w=0.7$ (blue) for $\omega=1$ and $\Delta=J=0$ i.e, qubits solely interact with the common environment not with each other. System is initially described by the mixture (91).
@@
In the next step we analyse ESD for finite temperatures. This time the Lindblad equation (77) must be considered. This equation can still be solved analytically, from which the concurrence for $\Delta=J=0$ is obtained as follows:

### v2

@@ III The model, the master equation and its solution
Our model system is a Heisenberg chain with only two localised qubits having the spin-spin interaction $(J_{x}S_{x}^{(1)}S_{x}^{(2)}+J_{y}S_{y}^{(1)}S_{y}^{(2)}+J_{z}S_{z}^{(1)}S_{z}^{(2)})/\hbar$ and are in interaction with an applied magnetic field along $z$-direction. The Hamiltonian for the $XY(J_{z}=0)$ model then recasts,
$H$ $=$ $\frac{J}{\hbar}(S_{+}^{(1)}\otimes S_{-}^{(2)}+S_{-}^{(1)}\otimes S_{+}^{(2)})+\frac{\Delta}{\hbar}(S_{+}^{(1)}\otimes S_{+}^{(2)}+S_{-}^{(1)}\otimes S_{-}^{(2)})$ (45)
@@
where $S_{\pm}^{(i)}=S_{x}^{(i)}\pm iS_{y}^{(i)}$ are the spin raising (lowering) operators for the $i^{th}$ qubit and $\omega$ is proportional to the intensity of the applied magnetic field. Coupling between qubits have been denoted as $J=(J_{x}+J_{y})/4$ and $\Delta=(J_{x}-J_{y})/4$ where the latter quantifies the anisotropicity of the interaction. It is noteworthy that systems with two coupled qubits have been utilized extensively in various contexts such as cavity QED [62] and nuclear spin systems with the hope of possible applications in quantum information schemes and quantum computing [1, 2].
At zero temperature, $T=0$, the state of the system evolves under the Lindbladian master equation [48]
$\dot{\rho}(t)$ $=$ $\frac{-i}{\hbar}[H,\rho(t)]+\frac{\gamma}{\hbar^{2}}~{}\mathcal{D}[S_{-}^{(1)}\otimes\mathds{1}^{(2)}]\rho(t)+\frac{\gamma}{\hbar^{2}}~{}\mathcal{D}[\mathds{1}^{(1)}\otimes S_{-}^{(2)}]\rho(t),$ (46)
where $\mathcal{D}[A]B=ABA^{\dagger}-\{A^{\dagger}A,B\}/2$ is the Lindblad super-operator and $\{X,Y\}$ denotes the anti-commutator of operators $X$ and $Y$. In our open quantum system this Lindblad equation describes the population relaxation of the upper state of each qubit with the relaxation rate $\gamma$ which for simplicity it has been taken equal for both qubits. In other words, interaction of qubits with the environment is assumed to be the same. This assumption is unjustifiable unless the interactions between spins are significantly weaker in comparison to the interaction of spins with the applied magnetic field. In other words, to be a valid assumption, the inter-particle interactions should not significantly change separations between the energy levels. The single-particle energy levels are ±$\pm\hbar\omega/2$ while those of the Hamiltonian (45) are given by ±$\pm\hbar J$ and ±$\pm\hbar\Omega$ where $\Omega=\sqrt{\Delta^{2}+\omega^{2}}$. Thus, the above requirements imply $J/\omega\ll 1$ and $\Delta/\omega\ll 1$ or equivalently $(\Omega-\omega)/\omega\ll 1$. In our calculations, these ratios are taken of the order of $10^{-1}$, indicating that the relative variations in the separation between the energy levels are less than 10 %.
The Lindblad equation (46) maintains the shape X of the state i.e., if the initial state is a X-shaped i.e., of the form (2) then the evolved state $\rho(t)$ is too.
@@
where $|\Psi_{-}\rangle=\frac{1}{\sqrt{2}}(|01\rangle-|10\rangle)$ is one of the maximally entangled Bell states, and $p$ a real number with constraint $p\in[-\frac{1}{3},1]$ which comes from the non-negativity requirement of the density matrix. For $p\in[-\frac{1}{3},\frac{1}{3}]$ the Werner state is separable and for $p\in[-\frac{1}{3},\frac{1}{\sqrt{2}}]$ it is local [63]. From this one sees that the Werner state is entangled but local for $p\in(\frac{1}{3},\frac{1}{\sqrt{2}}]$ i.e., entanglement does not imply nonlocality. The purity of the Werner state (56) is $\text{tr}(\rho_{W}^{2})=(1+3p^{2})/4$ being one for $p=1$ and $1/4$ for $p=0$. The former implies the pure ensemble $|\Psi_{-}\rangle\langle\Psi_{-}|$ while the latter corresponds to the completely random ensemble. The solution of the Lindblad equation (46) this time is given by
@@
where $\Omega$ is given by (55). As one sees the state $\rho_{W}(t)$ is independent of the parameter $J$ appearing in the Hamiltonian being responsible for the isotropic part of the spin-spin interaction. An interesting point about the Werner state is that for $\gamma=0$, the density operator is independent of time, which is the manifestation of the fact that the Werner state (56) commutes with the Hamiltonian (45) of the system, $[H,\rho_{W}(t)]=0$.
@@
for the steady state to be entangled. As we saw, reduced density operators for a X-shaped state are both always diagonal with zero quantum coherence. However, Eq. (64) shows that the global steady state $\ell_{1}$-norm coherence is,
@@ IV Temperature dependence of the steady state solution
$\dot{\rho}(t)$ $=$ $\frac{-i}{\hbar}[H,\rho(t)]+\frac{\gamma}{\hbar^{2}}(\bar{n}+1)~{}\mathcal{D}[S_{-}^{(1)}\otimes\mathds{1}^{(2)}]\rho(t)+\frac{\gamma}{\hbar^{2}}(\bar{n}+1)~{}\mathcal{D}[\mathds{1}^{(1)}\otimes S_{-}^{(2)}]\rho(t)$ (67)
$+\frac{\gamma}{\hbar^{2}}\bar{n}~{}\mathcal{D}[S_{+}^{(1)}\otimes\mathds{1}^{(2)}]\rho(t)+\frac{\gamma}{\hbar^{2}}\bar{n}~{}\mathcal{D}[\mathds{1}^{(1)}\otimes S_{+}^{(2)}]\rho(t),$
$\bar{n}$ being the average excitation of the thermal bath parametrizing the temperature: being zero for $T=0$ [48]. This equation is known as quantum optical master equation [64]222page: 148, equation (3.219) where the first two terms of the non-unitary part in the right side of the equation, describe the decay process while the remaining part i.e., the last two terms on the right hand side, are responsible to the excitation process with a rate that depends on temperature through $\bar{n}$. Note that in the limit of zero temperature, this equation reduces to Eq. (46) which describes the spontaneous decay, a purely dissipative process [65]. Eq. (67) can be analytically solved for the X-state (2). Again, one sees that this master equation preserves the shape of the state. Here, we are only interested in the steady state, $t\to\infty$, solution which is given by
@@
respectively for the steady state concurrence, CC and MIN. In this case, the $W$ matrix (31) is diagonal with $W_{11}(\infty)=W_{22}(\infty)$. Thus, for the steady state LQU one has that
@@
For non-interacting qubits, $J=\Delta=0$, the steady state density matrix recasts to the diagonal form,
$\rho(\infty)$ $=$ $\text{diag}\left\{\frac{\bar{n}^{2}}{(2\bar{n}+1)^{2}},\frac{\bar{n}(\bar{n}+1)}{(2\bar{n}+1)^{2}},\frac{\bar{n}(\bar{n}+1)}{(2\bar{n}+1)^{2}},\frac{(\bar{n}+1)^{2}}{(2\bar{n}+1)^{2}}\right\}$ (81)
implying that all coherences are zero. Then, from (19) one obtains
$\varrho_{ee}(\infty)=\frac{\bar{n}^{2}}{(2\bar{n}+1)^{2}},\quad\varrho_{gg}(\infty)=\frac{(\bar{n}+1)^{2}}{(2\bar{n}+1)^{2}},\quad\varrho_{ss}(\infty)=\varrho_{aa}(\infty)=\frac{\bar{n}(\bar{n}+1)}{(2\bar{n}+1)^{2}}\quad\varrho_{sa}(\infty)=\varrho_{eg}(\infty)=0$ (82)
for the elements of the density matrix in the Dicke basis. These relations reveal that all steady-state coherences vanish; and the symmetric and antisymmetric states are equally populated in the stationary regime. Then, from Eqs. (20), (21) and (4) it is found that for non-interacting qubits, the steady state concurrence is zero.
@@ V Results and discussion
In this section, we aim to consider the impact of the relaxation rate $\gamma$, responsible for the interaction of our qubit-qubit Heisenberg XY system with the environment, on the dynamics of quantum correlations by fixing other parameters of the problem including coupling strengths $J$ and $\Delta$; and magnetic field intensity $\omega$. However, as we saw in the previous section, Werner states and thus corresponding quantum correlations are independent of $J$. The same is valid for the general X-shaped state in the steady situation. Furthermore, in some calculations we fix all parameters other than $\Delta$ to study anisotropicity dependence of quantum correlations. Our investigation will primarily focus on examining how the zero-temperature steady state quantum correlations appear for different initial X states, including a mixture and the Werner state. After, the interesting phenomenon of entanglement sudden death, for independent qubits, will be considered for two initially entangled and afterwards not interacting qubits both for zero and finite temperatures. The time of ESD will be obtained analytically (numerically) for zero (finite) temperatures. Environment-induced quantum correlations is another phenomenon which will be considered. At the end finite temperature quantum correlations will be examined in the steady state case by studying their behaviour with temperature.
@@
Figure 1 depicts evolution of different quantum correlations for $J=0.1\omega$ and $\Delta=0.5\omega$ for three different values of the relaxation rate $\gamma=0.1\omega$, $\gamma=0.15\omega$ and $\gamma=0.2\omega$. Here, system is initially described by the entangled mixture (47). Initial values of the quantum correlations, being independent of $\gamma$, are $\mathcal{C}(0)=0.5$, $\mathcal{L_{N}}(0)=0.2716$, $\mathcal{U}(0)=0.5$ and $C_{cc}(0)=0.5$. As this figure reveals, both entanglement quantifiers, concurrence and log negativity, exhibit several dark and revival periods [19] before reaching the steady state. Because of the spontaneous emission, the initial entanglement diminishes with time, disappears in a given finite time which itself depends on the value of $\gamma$ and after a dark while it revives. The number and length of these dark intervals both depend on the value of $\gamma$. The first dark period is the longest one, which itself decreases with $\gamma$. The difference between entanglement quantifiers, $\mathcal{C}$ and $\mathcal{L_{N}}$, and quantum correlations beyond entanglement, $\mathcal{U}$ and $C_{cc}$, is that the former ones revive after zero entanglement, while the others can revive immediately several times without completely vanishing for a period of time; there are no dark periods for quantum correlations beyond QE. Moreover, LQU captures less quantum correlations than the other quantifiers, with the former converging to a stationary value of $0.1597$ for $\gamma=0.1\omega$. As one expects, the transition to the steady state occurs sooner for higher values of $\gamma$ and as one can check analytically from Eq. (64) the steady state value of quantum correlation although is independent of the initial state but depends on the value of $\gamma$.
@@
Figure 1: Evolution of the quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel), for $\gamma=0.1\omega$ (red curves), $\gamma=0.15\omega$ (green curves), $\gamma=0.2\omega$ (blue curves) for $J=0.1\omega$ and $\Delta=0.5\omega$. The quantum system is initially described by the mixture (47).
@@
Then one has that $\xi=\text{tr}(\rho_{\text{mix}}^{2})=1-2w(1-w)$ for the purity of the state. Time dependent density matrix obtained from the solution of the Lindblad equation (46) with $J=\Delta=0$ and the initial condition (83) is then,
@@
It follows that in the stationary regime, $t\to\infty$, all coherences vanish and all populations except $\rho_{44}$ become zero. This implies that, in this limit, the non-interacting qubits accumulate in the ground state as a result of spontaneous decay due to interaction with the environment. Note that, as we have shown in Eq. (64), this behaviour is independent of the initial state. Now, we consider evolution of concurrence. For this state, $\mathcal{C}_{2}(t)$ given by (6) is negative, having no contribution to the entanglement; only the two-photon coherence $\mathcal{C}_{1}(t)$ is present. Thus, from the definition (4) one obtains
@@
for the concurrence of this mixture. This equation shows that this state is always separable for $w=1$ as it is obvious from (83) for the initial state. For $10\leq w<$ 1 which the initial state is entangled, Eq. (91) shows that there is a finite time when the entanglement disappears, the signature of ESD. This time is obtained as
@@
for the time of ESD which is infinite for $w=0$ meaning that the entanglement does not die suddenly for the maximally entangled state $\frac{1}{2}(|00\rangle+|11\rangle)(\langle 00|+\langle 11|)$. The age of entanglement is shorter for higher values of both $\gamma$ and $w$. See figure 2. The weight factor $w$ is itself related to purity of the state (83) through $\xi=\text{tr}(\rho_{\text{mix}}^{2})=1-2w(1-w)$. As a result, by engineering the purity of the initial state, ESD can be slowed down.
The origin of ESD can be understood with regard to the populations and coherences; and the rates with which these quantities decay.We have considered non-interacting qubits for which $J=0=\Delta$. Thus, the Hamiltonian (45) recasts in $H=\omega(|00\rangle\langle 00|-|11\rangle\langle 11|)$ with eigenvalues $\{\omega,0,0,-\omega\}$ and corresponding eigenvectors $|00\rangle$, $|01\rangle$, $|10\rangle$ and $|11\rangle$. From the linear superpositions of $|01\rangle$ and $|10\rangle$, one then constructs symmetric and antisymmetric eigenvectors. In this way, four vectors of the Dicke basis (7), (8), (9) and (10) are obtained. ESD is then explained by analysing Eq. (20) in terms of the populations of the symmetric and antisymmetric states, $\varrho_{ss}(t)$ and $\varrho_{aa}(t)$, and the rates with which the populations and the two-photon coherence $|\varrho_{eg}(t)|$ decay [66].
Figure 2: Left panel: time of entanglement death $\gamma\tau$ given by Eq. (92) in terms of the weight $w$ appearing in the mixture (83). Right panel: evolution of concurrence (91) for $w=0.1$ (black curve), $w=0.3$ (red curve), $w=0.5$ (green curve) and $w=0.8$ (blue curve). Parameters have been chosen $\Delta=J=0$ i.e, qubits solely interact with the common environment, not with each other.
@@
for CC or equivalently for MIN with respect to (36). This equation shows that $C_{cc}(0)=1-w$, being nonzero for $w\neq 1$, and CC decreases exponentially; the usual behaviour of decaying systems. Thus, there is no phenomenon of CC sudden death, i.e., this quantity does not disappear in a finite time. For LQU which is computed by (33) analytical calculations show that the square root expression appearing in this equation is zero. Thus, LQU for $J=\Delta=0$ is obtained from
@@
where $W_{ij}$ is given by (31). The initial value of LQU is $\mathcal{U}(0)=\max\{1-w,1-\sqrt{w(1-w)}\}$ being zero for $w=1$. In the left panel of Fig. 3 we have plotted initial LQU versus the probability weight. As the numerical computations in the right panel of Fig. 3 show, LQU decreases first reaching a local minimum then increases to a given value and breaks downwards and continuously, not abruptly, vanishes. Ref. [68] has already reported this sudden change in LQU when one of the qubits is subjected to local decoherence noise. In any way, there is no LQU sudden death. Note that breaking points in the curves are due to the “max” operation in the definition (94).
@@
Figure 3: Initial value of LQU in terms of the probability weight $w$ (left panel) and evolution of LQU (right panel) for $w=0.4$ (cyan), $w=0.5$ (red), $w=0.6$ (green) and $w=0.7$ (blue) for $\Delta=J=0$ i.e, qubits solely interact with the common environment not with each other. The system is initially described by the mixture (83).
Next, we illustrate in figure 4 the dynamics of quantum correlations for the Werner state (56) for different values of $p$. Initial values of quantum correlations are $\mathcal{C}(0)=\max\{0,|p|-\frac{1-p}{2}\}$, $\mathcal{L_{N}}(0)=\log_{2}\left(\frac{1+p}{2}+\frac{1}{2}\left|\frac{1-p}{2}-|p|\right|+\frac{1}{2}\left|\frac{1-p}{2}+|p|\right|\right)$, $C_{cc}(0)=|p|$ and $\mathcal{U}(0)=1-\frac{1}{2}[1-p+\sqrt{(1-p)(1+3p)}]$ from which one finds the concurrence is non-zero only for $\frac{1}{3}<p\leq 1$ implying, as is well-known, that in this region the Werner state (56) is entangled. While QE, quantified by both concurrence and log negativity, is initially zero for values $p=-\frac{1}{3}$ and $p=0$ after a while it is born and finally reaches the steady value $0.2999$ and $0.3784$, being independent of $p$, respectively for concurrence and log negativity for $\gamma=0.1\omega$ and $\Delta=0.5\omega$. Using Eq. (64) one obtains $\mathcal{C}(\infty)=2\max\left\{0,\frac{|\Delta|\sqrt{4\omega^{2}+\gamma^{2}}-\Delta^{2}}{4\Omega^{2}+\gamma^{2}}\right\}$ for the steady state value of quantum concurrence which is non-zero only for $|\Delta|<\sqrt{4\omega^{2}+\gamma^{2}}$ As pointed out in the previous section, for $\gamma=0$, the Werner state (56) does not evolve in time, meaning that all quantum correlations remain constant. From all these we see that if both conditions $\frac{1}{3}<p\leq 1$ and $|\Delta|<\sqrt{4\omega^{2}+\gamma^{2}}$ fulfil simultaneously then for $|p|-\frac{1-p}{2}<\frac{|\Delta|\sqrt{4\omega^{2}+\gamma^{2}}-\Delta^{2}}{4\Omega^{2}+\gamma^{2}}$, the steady state is more entangled than the initial state. For parameters $\gamma=0.1\omega$ and $\Delta=0.5\omega$ where $\mathcal{C}(\infty)=0.2999$, this condition fulfils for $\frac{1}{3}<p<0.5333$.
Figure 4: Evolution of the quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel), for $p=-\frac{1}{3}$ (black curves), $p=0$ (red curves), $p=\frac{1}{\sqrt{3}}$ (green curves), and $p=1$ (blue curves) for $\gamma=0.1\omega$ and $\Delta=0.5\omega$. Note that the Werner state is independent of the isotropic part of qubits’ interaction, i.e., the coefficient $J$.
Figure 5: Evolution of the quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel) for different values of the anisotropy interaction coefficient: $\Delta=0$ (brown), $\Delta=0.2\omega$ (magenta) $\Delta=0.3\omega$ (cyan) and $\Delta=0.4\omega$ (orange); and for $\gamma=0.1\omega$. System is initially described by the Werner state (56) with $p=0$ (completely random ensemble; $\rho_{0}=\frac{1}{4}\mathds{1}_{4}$).
After, we highlight the constructive role played by the environment in favour of quantum correlations; the so-called environment-induced or decoherence-induced quantum correlations. To this end, we consider our system is initially described by the Werner state (56) with $p=0$ i.e., $\rho_{0}=\frac{1}{4}\mathds{1}_{4}$; the completely random ensemble which contains no correlations. Then, evolution of quantum correlations is considered for different values of the interaction strength $\Delta$. Note that the evolved state and thus quantum correlations are all independent of the value of $J$. Figure 5 depicts variation of quantum correlations with time for a given value of $\gamma$ but different values of interaction strength $\Delta$. For $\Delta=0$ all quantum correlations are always zero but for $\Delta\neq 0$ quantum correlations are generated as time elapses. From this point we learn that (i) environment plays a constructive role, in contrast to its customary destructive role, inducing quantum correlations (ii) in addition to the interaction of qubits with the environment, the anisotropic part of their own interaction with each other is necessary for creation of quantum correlations. Generally, to see the possible creation/enhancement of a specific quantum correlation $\mathcal{Q}$, one should compare its steady state value with its initial value. In this way, the condition $\mathcal{Q}(\rho_{\infty})>\mathcal{Q}(\rho_{0})$ implies the phenomenon of creation/enhancement of the quantum correlation $\mathcal{Q}$. As top plots of this figure show, generation of QE from an initially separable state, the so-called environment-induced entanglement, requires both interaction between qubits and also their own interaction with the environment; no one is sufficient on its own. This constructive role played by the environment in favour of entanglement, in contrast to its customary destructive role, has already been noticed. See for instance [69, 70]. Despite the apparent implication from this figure that quantum correlations increase with the strength of interaction $\Delta$, it is important to note that this is not always the case. Consider the steady state as a example. For our parameters, $\gamma=0.1\omega$ the steady state concurrence $\mathcal{C}(\infty)=2\max\left\{0,\frac{|\Delta|\sqrt{4\omega^{2}+\gamma^{2}}-\Delta^{2}}{4\Omega^{2}+\gamma^{2}}\right\}$ is nonzero only for $\frac{\Delta}{\omega}\in(-2.0025,0)\cup(0,2.0025)$ and is an increasing function of $\Delta$ only when $-2.0025~{}\omega<\Delta\leq-0.6188~{}\omega$ and $0<\Delta\leq 0.6188~{}\omega$.
@@
In the next step we analyse ESD for finite temperatures. This time, the Lindblad equation (67) must be considered. This equation can still be solved analytically, from which the concurrence for $\Delta=J=0$ is obtained as follows: